OpenStudy (anonymous):

find the standard form of the complex # 8(cos(30degrees)+i sin(30degrees))

OpenStudy (anonymous):

you know what \(\sin(30)\) and \(\cos(30)\) are ? if so, make the replacements and get a number

OpenStudy (anonymous):

doesnt make sense

OpenStudy (mertsj):

\[8(\frac{\sqrt{3}}{2}+i\frac{1}{2})=4\sqrt{3}+4i\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw